Use the poisson formula to find the probability of the value given for the random variable x.

1. M = 2, x = 3           2. M = 4, x = 1          3. M = 0.845, x = 2    4. M = 0.250, x = 2
Answer #1

Solution:

1)Given that ,

mean = M = 2

Using Poisson probability formula,

P(X = x) = (e-M * Mx ) / x!

P(X = 3) = (e-2* 2 3) / 3! = 0.18045

2) Given that ,

mean = M = 4

Using Poisson probability formula,

P(X = x) = (e-M * Mx ) / x!

P(X = 1) = (e-4* 41) / 1! = 0.07326

3)Given that ,

mean = M = 0.845

Using Poisson probability formula,

P(X = x) = (e-M * Mx ) / x!

P(X = 2) = (e-0.845* 0.8452) / 2! = 0.15336

4) Given that ,

mean = M = 0.250

Using Poisson probability formula,

P(X = x) = (e-M * Mx ) / x!

P(X = 2) = (e-0.250* 0.2502) / 2! = 0.02434